OpenStudy (anonymous):

x + y = 6 2x - y = 3 The system of equations has no solution has one solution is coincident

OpenStudy (anonymous):

well the fastest approach is to sum the equations: x+y=6 + 2x-y=3 = 3x=9 so x=3 so we get 3+y=6 y=3 as well so the system has one solution which is x=3, y=3